How much energy is required to heat 1.5 L of water at a temperature of 20 °C to the boiling point 100 ºC?

Answer by rfadrogane(214)

How much energy is required to heat 1.5 L of water at a temperature of 20 °C to the boiling point 100 ºC?
Sol'n:
Q = mCp(T2-T1)
m = mass density of water = 1 kg/L for Water, thus m = 1.5-kg
Cp = 4.186 kJ/kg-ºC, T2 = 100ºC & T1 = 20ºC
Then,
Q = 1.5(4.186)(100-20)
Q = 502.32 kJ ----answer
There is 1.5 L of water in a kettle on a cooking plate of 2.0 KW. It takes 7 minutes to heat the water from 20 C to the poiling point (100 C). How much electrical energy is required ? and what is the efficiency of the heating process?
Sol'n:
From above: Q = 502.32-kJ, energy needed to raise the temp. of water.
From: Q = (eff.)E
where: E - the energy needed by the kettle
eff.= efficiency of the kettle
E = Pt
= 2,0000 J/s x 7(60 s.)
E = 840-kJ ----answer
thus,
eff. = Q/E (100)
= 502.32/840 (100)
eff. = 59.8% ----answer
just send your question to my email:
rfadrogane@rocketmail.com